OpenStudy (kayy):

Wesley works at a shoe store. He is paid $8 an hour plus a 7% commission on his total sales. Which of the following best represents t, Wesley’s total weekly earnings if he works 30 hours and has total sales of s dollars?

OpenStudy (anonymous):

get out of here

OpenStudy (kayy):

No!! I need answers...I mean help(:

OpenStudy (anonymous):

i think u forgot to give the options becoz u say "which of the following"

OpenStudy (kayy):

Oops Sorry!(: F t = 0.08(30) + 0. G t = 8(30) + 7s H t = 8(30) + 0.07s J t = 8(30) + 0.7s

OpenStudy (kayy):

haha cute smiley face(:

OpenStudy (kayy):

Hahahaha :P

OpenStudy (anonymous):

Your answer is H (t) = 8(30) + 0.07s becoz he gets $8/h , so for 30 hrs he will get 8(30) also 7% = 7/100 = 0.07 so 7% of s = 0.07s hence total income = 8(30) + 0.07s

OpenStudy (kayy):

Thanks Harkirat(:

OpenStudy (anonymous):

welcome :))

OpenStudy (kayy):

(: